# SandCell: Sandboxing Rust Beyond Unsafe Code

### v1

@@ 5.3 Analysis and Instrumentation
After the user specifies the sandbox boundaries, the next step is to insert necessary code for enforcing sandboxes and passing data between sandboxes. Given the source code, SandCell builds a call graph of the program reusing Rupta [46], a Rust pointer analysis tool. Based on the specification and the call graph, SandCell identifies the public interfaces of types, modules and crates to be sandboxed. Then it inserts calls to the low-level isolation APIs at these interfaces to enforce the isolation at runtime.
@@
For low-level isolation, SandCell adopts an open-sourced isolation library SDRaD by Gülmez et al [35], and augments it as SDRaD-v2, adding syscall filtering based on zpoline [65] and mimalloc [45] as the memory allocator. More details can be found in §˜7.3. At a high level, SDRaD-v2 provides a set of APIs for managing sandboxes. To switch between sandboxes, SandCell inserts calls to these APIs at the entry and exit points of the sandboxed functions. Also, it provides APIs for management of share data domains for allocation of objects that two sandboxes can access. More details on the SDRaD-v2 APIs can be found in §˜7.
@@
The data passed across sandbox boundaries includes the arguments passed to the sandboxed functions and their return values. To make data from one sandbox available to another sandbox, one approach is to deep copy all such data as in Glamdring [47], PtrSplit [49], and GOTEE [33]. This approach, however, can lead to a high overhead since deep copying copies all transitively reachable data; it should be avoided when the size of transmitted data is large [34].
@@
Instead of deep copying, our design copies only stack data and allocate heap data (when it does not contain references to stack data) on a shared data domain that is accessible from both sides of a cross-boundary call. This approach reduces the overhead of copying cross-boundary data. On the other hand, the complexity is that cross-boundary data may not be allocated next to each other so that we can easily make relevant memory pages a shared memory region. If we have to copy such data to a shared data domain during cross-boundary calls, we are back to the same problem of high copying cost. To address this, we propose a novel approach to (1) employ static analysis to compute allocation sites that may allocate cross-boundary data in memory, and (2) instrument those allocation sites so that the data is allocated in a shared data domain. This approach avoids the data copying cost by directly allocating cross-boundary data in a shared domain. We detail the static analysis in §˜6 and the instrumentation in §˜7.
5.4 Run-time Architecture
Figure 3: SandCell run-time architecture.
@@ 5.4 Run-time Architecture
Figure˜3 shows an overview of the SandCell architecture at runtime. When the program starts, SDRaD-v2 places the main function and the program stack and heap into a special root domain (➀). When SandCell creates a sandbox (➁), a new SDRaD-v2 domain (➂) is initialized and a new, dedicated, stack and heap area is allocated. Code in the SDRaD-v2 root domain can access data in any new domain created by SandCell, but sandboxed code cannot access the stack or heap of other domains, including the root domain. SDRaD-v2 control data is stored with a separate, monitor domain (➃) that is inaccessible to code in the root or sandboxed domains. The monitor domain will run the code for domain management (➄), the domain-aware memory allocator (➅), and the system call filtering (➆). We will discuss the low-level details in §˜7.
@@
Notice that, as shown in Figure˜3, crate C is a dependency of both crate A and crate B. However, at runtime, because crate A and crate B are sandboxed separately, each of them has its own copy of the code and data of crate C.
5.5 Security Guarantee
@@
SandCell provides memory isolation between sandboxes, preventing memory vulnerabilities from being exploited to access memory outside the sandbox. For illustration, suppose that in Figure˜3, the attacker exploits a memory vulnerability in crate A and gains the ability to read and write arbitrary memory. Because of the memory isolation provided by SandCell, the attacker can only directly manipulate the memory within Domain #1. Attempts to access memory in other domains, including the root domain, will result in a runtime error. As a practical example, CVE-2018-25032 [27] is a out-of-bound write triggered by crafted user input in zlib. If sandboxed by SandCell, the out-of-bound write can only modify memory within the sandbox. Another example, CVE-2023-0216 [29], is an invalid pointer dereferencing vulnerability in openssl. Sandboxing the library confines the invalid memory access within the sandbox.
@@
We note that the guarantee provided by SandCell is typical for sandboxing tools and it does not provide interface safety [44], which requires the data and control flow at the sandbox boundaries be carefully validated to prevent potential misuses or abuse. How to provide interface safety effectively and efficiently in sandboxing is still an area of research.
6 Dataflow Analysis
This section is dedicated to the static analysis that computes allocation sites of cross-boundary data. We first define the problem and then describe the analysis algorithm.
6.1 Problem definition
At the call site of a sandboxed function, we need to determine whether for each argument (and every part of the argument, e.g., a field of a struct) can point to some heap object and, if so, compute a set of allocation sites where the heap object can be allocated. Once we know those allocation sites, in a later step we can instrument those sites so that they allocate data in a specialized data domain. For example, suppose we have a sandboxed function foo(x: Vec<i32>), and a call site foo(v); the analysis should determine an allocation site (or a set of sites) for v that performs an operation such as v = Vec::new().
We clarify how we identify what heap objects are: in Rust, heap objects are created by calling a memory allocator to reserve a piece of memory on the heap, similar to C. But different from C, this call necessarily goes through some unsafe code. Since using a large amount of unsafe code or scattering unsafe code blocks everywhere is highly discouraged in Rust, it is a common practice to wrap heap allocations in abstractions that expose safe APIs to the user. In fact, most heap objects used in practice are in the form of standard library containers, such as Vec, HashMap, etc. With this observation, in our analysis, we determine if a variable is a heap object by checking if its type is in a list of standard library containers. Users can also customize this list by including type names of their desired heap objects.
6.2 Static analysis
Type $\quad\tau$ $::=$ $\text{bool}\mid\text{i32}\mid\text{Vec}\langle\tau\rangle\mid\&\tau\mid\&\text{mut}\ \tau\mid\dots$
Local $\quad v$ $\in\ \$$\{v_{0},v_{1},v_{2},\dots,v_{m}\}\cup\{v_{m+1},\dots\}$
Function $\quad f$ $::=$ $\text{f}(v_{1},v_{2},\dots,v_{m})\;\{s;\}$
Statement $\quad s$ $::=$ $s_{1};s_{2}\mid e=r\mid e=f(e_{1},e_{2},\dots)\mid\dots$
Place $\quad e$ $::=$ $v\mid*e\mid e.\eta$
Rvalue $\quad r$ $::=$ $e\mid\&e$
Projection $\quad\eta$ $::=$ $\epsilon\mid n.\eta$
Figure 4: A simplified grammar of Rust MIR.
$e_{0}=\text{sandboxed}(\dots,e,\dots)\in\mathcal{S}$     $e.\eta\in\mathcal{P}$            $e.\eta\in\mathcal{R}$
$p_{1}=p_{2}\in\mathcal{S}$     $p_{2}.\eta\in\mathcal{R}$        $p1.\eta\in\mathcal{R}$
$p_{1}=\&p_{2}\in\mathcal{S}$     $p_{2}.\eta\in\mathcal{R}$        $p1.\eta\in\mathcal{R}$
$p=(*v).\eta_{0}\in\mathcal{S}$     $v.\eta_{0}.\eta\in\mathcal{R}$         $p.\eta\in\mathcal{R}$
$(*v).\eta_{0}=p\in\mathcal{S}$     $p.\eta\in\mathcal{R}$         $v.\eta_{0}.\eta\in\mathcal{R}$
$p=\&((*v).\eta_{0})\in\mathcal{S}$     $v.\eta_{0}.\eta\in\mathcal{R}$            $p.\eta\in\mathcal{R}$
     $e=f(e_{1},e_{2},\dots)\in\mathcal{S}$   $\{e=f.v_{0},f.v_{1}=e_{1},f.v_{2}=e_{2},\dots\}\subset\mathcal{S}$
Figure 5: Rules for computing the backward reachability set $\mathcal{R}$.
To compute allocation sites from arguments of calls to sandboxed functions, our static analysis performs backward analysis on a dataflow graph, which is a specialized form of aliasing analysis via backward reachability. The dataflow graph is an abstraction of the input program based on the program’s Rust Mid-level Intermediate Representation (MIR), which we obtain by running the Rust compiler.
We next formalize the static analysis using a set of rules on a simplified version of MIR proposed by Rupta [46], as shown in Figure˜4. Local variable $v_{0}$ is the return value of a function, and $v_{1},v_{2},\ldots,v_{m}$ are its arguments. Other locals $v_{m+1},\dots$ include variables in the source code and temporary variables generated by the compiler. Projection $\eta$ is a way to access a field of a struct, where $n$ is the field name. A place can refer to memory locations both on the stack and the heap. We also mark a subset of places as local places $p::=v.\eta$, which are places that have no dereference operator and can only refer to stack objects. Since our analysis is flow-insensitive, we omit some control flow statements in Figure˜4.
Figure˜5 shows a set of rules to compute places that might flow to arguments of sandboxed function calls. We denote the set of statements in the entire input program as $\mathcal{S}$, the set of all places as $\mathcal{P}$, and the backward reachable set of places as $\mathcal{R}$, which is computed by the rules. We also denote the return value of a function $f$ as $f.v_{0}$ and its arguments as $f.v_{1},f.v_{2},\ldots,f.v_{m}$. For simplicity, a place $(*p).\eta$ is denoted as $p.\eta$. Initially, by the first rule in Figure˜5, $\mathcal{R}$ includes all places that are arguments of sandboxed function calls and their fields. For each statement $s\in\mathcal{S}$, we update $\mathcal{R}$ according to the rules. For example, for the statement $p_{1}=p_{2}$, if a place $p_{2}.\eta$ is in $\mathcal{R}$, we add $p_{1}.\eta$ to $\mathcal{R}$. Other rules are similar. The last rule makes the analysis interprocedural by adding assignment statements to model parameter passing in a function call. We keep updating $\mathcal{R}$ and $\mathcal{S}$ until they reach a fixed point.
Given an $\mathcal{R}$ computed by the rules, we scan the program for all allocation function calls $v=\text{alloc}()$ that have $v\in\mathcal{R}$. These calls are the sites that can allocate cross-boundary data. Then we can change the allocation domain of those data by redirecting the allocation function calls to a shared domain.
An example
. Consider the following code snippet:
struct S { s: Vec<i32> }
@@ 6.2 Static analysis
let x = S { s: Vec::new() };
let y = Vec::new();
let z = if (/* condition */) {
&x.s
} else {
&y
sandbox(*z);
Suppose sandbox is a sandboxed function. The analysis notices that z used in a sandoxed function call is of type &Vec<i32>, which is a reference to a heap object. By the analysis, initially $\mathcal{R}$ includes z. Because of the branch statement, $\mathcal{R}$ is updated to include &x.s and &y, which further leads to the inclusion of x.s and y, respectively. Then the analysis scans the program and finds the two allocation sites of Vec::new() that are associated with x.s and y, which are possible allocation sites for the data z references. This example also shows that our static analysis is an over-approximation: even though exactly one of the two allocated vectors is crossing the boundary during execution, the analysis would conservatively return both allocation sites, resulting in both objects on the shared domain.
7 Implementation
@@ 7.1 Program Analysis
To implement the analysis discussed in §˜6 at the MIR level, we first construct a call graph of the program, reusing call graph construction of Rupta [46]. Rupta mimics the monomorphization process of the Rust compiler to resolve polymorphic functions to concrete ones and constructs a call graph between concrete functions.
Based on the call graph, we construct a dataflow graph. Each function in MIR has a list of Locals, including the function arguments, temporaries, local variables, and the return value. Each Local might contain a list of Places that correspond to memory locations, e.g., fields of a struct. We map each Place to a node in the dataflow graph and create edges between Places on the left-hand side and right-hand side of an assignment as discussed in §˜6. For MIR statements that do not fall into the categories listed in §˜6, we reuse Rupta’s statement transformation to convert them. Given the constructed dataflow graph, the algorithm in Figure˜5 is straightforward to implement. To reduce the false positive rate, we add context-sensitivity to the dataflow analysis by recording the call and return functions along the dataflow path and filtering out the paths that break the call-and-return matching relation.
7.2 Instrumentation
The instrumentation phase inserts code to enforce the sandbox boundaries specified by the user. After the analyzer processes the MIR of the entire program (including dependencies) and persists the result to disk, the instrumentor reads the analysis result and rebuilds the program.
For each function that crosses the sandbox boundary, a wrapper function is inserted to switch the context to the callee’s sandbox by calling SDRaD-v2 APIs, which will be elaborated in §˜7.3. Then the instrumentor redirects the original function call to the wrapper function.
For sharing heap objects, the instrumentor modifies the allocator of the heap object using the standard library’s allocator API. In Rust’s standard library, heap objects are generic over allocators; e.g., Vec<T, A: Allocator> can be instantiated with different allocators. By using Vec<i32, SandCellAlloc>, we direct all memory allocations associated with Vec to our custom allocator SandCellAlloc, which allocates memory in a shared data domain.
@@ 7.2 Instrumentation
7.3 In-Process Isolation
The low-level isolation mechanism used by SandCell is responsible for enforcing the isolation between different sandboxes and providing primitives to transition securely from one sandbox to another. SandCell implements hardware-assisted in-process isolation via the x86-64 processor PKU feature by building on an open-source isolation library by Gülmez et al [35], Secure Rewind & Discard (SDRaD), that implements lightweight primitives for in-process isolation and fault recovery. We augment SDRaD with support for a Rust-compatible memory allocator and harden it by incorporating a system call filtering engine. For clarity, we refer in the following description to SDRaD when functionality is provided by the original SDRaD library, and to SDRaD-v2 when describing added functionality.

### v2

@@ 5.3. Analysis and Instrumentation
After the user specifies the sandbox boundaries, the next step is to insert necessary code for enforcing sandboxes and passing data between sandboxes. Given the source code, SandCell builds a call graph and a data flow graph of the program. Based on the specification and the call graph, SandCell identifies the public interfaces of types, modules and crates to be sandboxed. Then it inserts calls to the low-level isolation APIs at these interfaces to enforce the isolation at runtime.
@@
For low-level isolation, SandCell adopts an open-sourced isolation library SDRaD by Gülmez et al (gulmez2023dsn, ), and augments it as SDRaD-v2, adding syscall filtering based on zpoline (yasukata23, ) and mimalloc (Leijen2019MimallocFL, ) as the memory allocator. More details can be found in § 7.3. At a high level, SDRaD-v2 provides a set of APIs for managing sandboxes. To switch between sandboxes, SandCell inserts calls to these APIs at the entry and exit points of the sandboxed functions. Also, it provides APIs for management of shared data domains for allocation of objects that two sandboxes can access. More details on the SDRaD-v2 APIs can be found in § 7.
@@
The data passed across sandbox boundaries includes the arguments passed to the sandboxed functions and their return values. To make data from one sandbox available to another sandbox, one approach is to deep copy all such data as in Glamdring (lind2017glamdring, ), PtrSplit (liu2017ptrsplit, ), and GOTEE (ghosn2019secured, ). This approach, however, can lead to a high overhead since deep copying copies all transitively reachable data; it should be avoided when the size of transmitted data is large (gulmezFriendFoeExploring2023, ).
@@
Instead of deep copying, our design copies only stack data and allocate heap data (when it does not contain references to stack data) on a shared data domain that is accessible from both sides of a cross-boundary call. This approach reduces the overhead of copying cross-boundary data. On the other hand, it also requires that cross-boundary data may not be allocated next to each other the relevant memory pages can be assigned to a shared memory region. If we have to copy such data to a shared data domain during cross-boundary calls, we are back to the same problem of high copying cost. To address this, we propose a novel approach to (1) employ static analysis to compute allocation sites that may allocate cross-boundary data in memory, and (2) instrument those allocation sites so that the data is allocated in a shared data domain. This approach avoids the data copying cost by directly allocating cross-boundary data in a shared domain. We detail the static analysis in § 6 and the instrumentation in § 7.
5.4. Run-time Architecture
Figure 3. SandCell run-time architecture.
@@ 5.4. Run-time Architecture
Figure 3 shows an overview of the SandCell architecture at run time. When the program starts, SDRaD-v2 places the main function and the program stack and heap into a special root domain (➀). When SandCell creates a sandbox (➁), a new SDRaD-v2 domain (➂) is initialized and a new, dedicated, stack and heap area is allocated. Code in the SDRaD-v2 root domain can access data in any new domain created by SandCell, but sandboxed code cannot access the stack or heap of other domains, including the root domain. SDRaD-v2 control data is stored with a separate, monitor domain (➃) that is inaccessible to code in the root or sandboxed domains. The monitor domain will run the code for domain management (➄), the domain-aware memory allocator (➅), and the system call filtering (➆). We will discuss the low-level details in § 7.
@@
Notice that, as shown in Figure 3, crate C is a dependency of both crate A and crate B. However, at run time, because crate A and crate B are sandboxed separately, each of them has its own copy of the code and data of crate C.
5.5. Security Guarantees
@@
SandCell provides memory isolation between sandboxes, preventing memory vulnerabilities from being exploited to access memory outside the sandbox. Suppose that in Figure 3, the attacker exploits a memory vulnerability in crate A and gains the ability to read and write arbitrary memory. Because of the memory isolation provided by SandCell, the attacker can only directly manipulate the memory within Domain #1. Attempts to access memory in other domains, including the root domain, will result in a runtime error. As a practical example, CVE-2018-25032 (CVE-2018-25032, ) is a out-of-bound write triggered by crafted user input in zlib. If sandboxed by SandCell, the out-of-bound write can only modify memory within the sandbox. Another example, CVE-2023-0216 (CVE-2023-0216, ), is an invalid pointer dereferencing vulnerability in openssl. Sandboxing the library confines the invalid memory access within the sandbox.
@@
We note that the guarantee provided by SandCell is typical for sandboxing tools and it does not provide interface safety (SandboxingSok2025SP, ), which requires the data and control flow at the sandbox boundaries be carefully validated to prevent potential misuses or abuse. How to provide interface safety effectively and efficiently in sandboxing is still an area of research.
6. Dataflow Analysis
We next describe the static analysis that identifies which heap objects may cross sandbox boundaries. We first illustrate the problem with a motivating example, then describe our analysis algorithm.
6.1. Problem definition
Goal.
When a sandboxed function is called, any heap data passed as arguments must reside in a memory region accessible to both the sandbox and the main program. Our analysis identifies these cross-boundary heap objects at compile time, so their allocation sites can be instrumented to use the shared memory domain.
Motivating example.
Consider a common pattern where data is accumulated through helper functions, illustrated by the following example, where the main function first loads a series of plugins and then invokes each plugin through a sandboxed call.
fn load_plugins() -> Vec<Plugin> {
let mut plugins = Vec::new(); // Allocation
for path in discover_paths() {
if let Some(p) = parse_plugin(path) {
plugins.push(p); // Container insert
plugins
fn run_sandbox(plugins: &[Plugin]) {
for p in plugins.iter() {
sandbox(p); // Sandboxed call
@@ 6.1. Problem definition
let plugins = load_plugins();
run_sandbox(&plugins);
The analysis must determine that allocation in load_plugins flows into the sandboxed call in run_sandbox. This requires tracking data through: (1) a return value flowing from load_plugins to main; (2) a function argument flowing from main into run_sandbox; (3) an iterator plugins.iter() that borrows the container; and (4) the push operation that inserts elements into the vector.
Without interprocedural analysis, we cannot connect allocations in load_plugins to uses in run_sandbox. Without modeling container operations, we cannot determine that data pushed into plugins is accessible through iteration. Our analysis handles these cases by building a dataflow graph that is interprocedural and includes summary edges for standard library container methods.
Identifying heap objects.
In Rust, heap allocations are wrapped in safe abstractions like Vec, HashMap, Box, etc. Our analysis recognizes allocations by detecting constructor calls for these standard library containers (e.g., Vec::new(), HashMap::with_capacity()).
6.2. Analysis algorithm
Node Type Description
Local variable A variable declared in a function
Parameter A function’s input parameter
Return value A function’s return value
Field projection A struct field access (e.g., x.field)
Dereference projection A pointer dereference (e.g., *ptr)
Heap allocation A call to Vec::new(), Box::new(), etc.
Edge Type Created For
Assign Direct assignment: x = y
Ref Reference creation: x = &y
Aggregate Struct/tuple construction
CallArg Argument $\to$ parameter at call sites
Return Return value $\to$ destination at call sites
StdSummary Standard library function summaries
Figure 4. Node and edge types in the dataflow graph.
Our analysis operates on Rust’s Mid-level Intermediate Representation (MIR), which the compiler generates before optimization. We construct an interprocedural dataflow graph (DFG) that captures how values flow through the program, then perform backward reachability to find allocation sites.
Dataflow graph structure.
The DFG is a directed graph where:
  • Nodes represent program locations that hold values: local variables, function parameters, return values, struct fields, and heap allocation sites.
  • Edges represent data flow from sources to destinations. An edge from node $A$ to $B$ means the value at $A$ may flow to $B$.
Figure 4 summarizes the node and edge types in our DFG.
We create edges for the following program constructs:
  1. (1)
    Assignments (x = y): Add an Assign edge from y to x.
  2. (2)
    References (x = &y): Add a Ref edge from y to x. We also handle reborrowing patterns like x = &*y, where we first create a projection node for *y, then add a Ref edge to x.
  3. (3)
    Dereferences (x = *y): Add an Assign edge from the projection node *y to x. The dereference node is connected to the pointer’s base via an Aggregate edge.
  4. (4)
    Struct/tuple construction (s = S { f: v }): Add an Aggregate edge from v to the field node s.f, and another Aggregate edge from s.f to s.
  5. (5)
    Function calls (r = f(a, b)): Add CallArg edges from each argument to the corresponding parameter of f, and a Return edge from f’s return value to r.
  6. (6)
    Heap allocations (v = Vec::new()): Create a heap allocation node and add an Assign edge to v.
For standard library functions whose bodies are not available or too complex to analyze, we use pre-defined summaries to specify which arguments flow to the return value. For example, Vec::iter() returns an iterator over its receiver; so we add a StdSummary edge from the receiver to the result.
Handling indirect calls.
Function pointers and trait objects create indirect calls where the target is not statically known. The analysis tracks how function pointers flow through the DFG and uses this information to resolve possible call targets. Critically, DFG construction and call target resolution are performed iteratively: we first build an initial DFG using only direct calls, then resolve indirect call targets based on dataflow, add edges for newly discovered calls, and repeat until no new targets are found. This iterative refinement ensures that the call graph and DFG are mutually consistent.
Backward reachability.
Given the DFG, we find allocation sites by traversing edges backward (from destination to source) starting from sandboxed call arguments. The algorithm is a straightforward worklist-based graph traversal: we initialize the worklist with all arguments of sandboxed function calls, then repeatedly pop a node, follow its incoming edges, and add unvisited source nodes back to the worklist. When a heap allocation node is reached, we record it as a cross-boundary allocation. The traversal continues until the worklist is empty.
Context sensitivity.
The backward traversal is call-site sensitive: when traversing a Return edge from callee to caller, we record the call site on a stack. Later, when traversing a CallArg edge, we check that it matches the recorded call site. This prevents spurious paths that would “return” to a different caller than the one that made the call. The analysis uses a configurable stack depth limit to bound analysis time while maintaining precision for common patterns.
7. Implementation
@@ 7.1. Program Analysis
In MIR, each function has a list of Locals, including the function arguments, temporaries, local variables, and the return value. Each Local might contain a list of Places that correspond to memory locations, e.g., fields of a struct. We map each Place to a node in the dataflow graph and create edges between Places on the left-hand side and right-hand side of an assignment as discussed in § 6. The rest of analysis has been described in § 6.
7.2. Instrumentation
The instrumentation phase inserts code to enforce the sandbox boundaries specified by the user. For each function that crosses the sandbox boundary, a wrapper function is inserted to switch the context to the callee’s sandbox by calling SDRaD-v2 APIs, which will be elaborated in § 7.3. Then the instrumentor redirects the original function call to the wrapper function.
For sharing heap objects, the instrumentor modifies the allocator of the heap object using the standard library’s allocator API. In Rust’s standard library, heap objects are generic over allocators; e.g., Vec<T, A: Allocator> can be instantiated with different allocators. By using Vec<i32, SandCellAlloc>, we direct all memory allocations associated with Vec to our custom allocator SandCellAlloc, which allocates memory in a shared data domain. For types that do not support custom allocators, SandCell wraps their allocations and all the following mutations with a macro that redirects the allocations to the shared data domain.
@@ 7.2. Instrumentation
7.3. In-Process Isolation
The low-level isolation mechanism used by SandCell is responsible for enforcing the isolation between different sandboxes and providing primitives to transition securely from one sandbox to another. SandCell implements hardware-assisted in-process isolation via the x86-64 processor PKU feature by building on an open-source isolation library, Secure Rewind & Discard (SDRaD(gulmez2023dsn, ), that implements lightweight primitives for in-process isolation and fault recovery. We augment SDRaD with support for a Rust-compatible memory allocator and harden it by incorporating a system call filtering engine. For clarity, we refer in the following description to SDRaD when functionality is provided by the original SDRaD library, and to SDRaD-v2 when describing added functionality.